OpenStudy (anonymous):

A right triangle has base (x - 5) units and height (4x - 10) units. Part A: What is the square of the length of the hypotenuse of the triangle? Show your work. (4 points) Part B: What is the area of the triangle? (3 points) Part C: Using the solution obtained in Part B, explain the closure property of multiplication of polynomials. (3 points) HELPPPP PLEASSEEEEEEE NEEDDD TO GET THIS DONEEEE WILL GIVE A MEDAL AND EVERYTHING!!!!!!

OpenStudy (anonymous):

hypotenuse^2 = base^2 + height^2

OpenStudy (anonymous):

Use the Pythagorean theorem to find the hypotenuse. You know the two sides. a = x-5 b = 4x-10 c is hypotenuse a^2 + b^2 = c^2 So square a, square b, add them together and take the square root. The result is c, the length of the hypotenuse. To find the area of a triangle, multiply 1/2 times the base times the height.

OpenStudy (anonymous):

area = (1/2) base x height